normal distribution with a mean of 61 and standard deviation of 5, how to calculate the standard score equivalents for
x x z t
66

Answer by stanbon(75887)

normal distribution with a mean of 61 and standard deviation of 5, how to calculate the standard score equivalents for
x x z t
66
-----
Not sure what you mean be repeating the x value.
--------
z(66) = (66-61)/5 = 5/5 = 1
-----
t(66) = is the same as z(66)
====================
Cheers,
Stan H.